Thirty-four

different video games showing

violence

were observed. The duration times of

violence

​(in seconds) were recorded. When using this sample for a t test of the claim that the population mean is greater than

9595 sec, what does df​ denote, and what is its​ value?
Answer #1

here df represent the number of sample values one can change without changing the estimate value

therefore degree of freedom =sample size(n)-1 =34-1 =33